OpenStudy (anonymous):

Help ; 5(3x - 2) - 4(3x - 4) = 0

OpenStudy (anonymous):

First you have to substitute you values.

OpenStudy (anonymous):

15x-10-12x+16=0 3x=10-16 3x=-6 x=-2

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

Once you substitute you should have 15x-10-12x+16=0 Then combine like terms 3x+6=0 3x=-6 x=-2